# A Hybrid ABM-PDE Framework for Real-World Infectious Disease Simulations

### v1

@@ 6 Discussion
The deterministic PDE dynamics and the coupling process both reduce variability in the 25% scenario, but this effect does not hold for the full population. In fact, the hybrid model introduces a different type of noise due to the abstraction and simplification of agent-level dynamics. Agent identifiers are lost in the PDE region, leading to mismatches when assigning states upon re-entry to the ABM. For instance, an agent may reappear earlier or later than expected, or with an incorrect health status. To mitigate this, we propose future improvements such as storing metadata when an agent transitions into the PDE domain – including expected return time and current health status. Upon re-entry, this information could be used to assign a health state based on time spent in the PDE region and the transition dynamics from the compartment model (see Fig. 1). This strategy would help reduce artificial variance introduced by random reassignment. Additionally, stochasticity could be reintroduced into the PDE model itself – for example, through noise terms in the compartment equations – to better reflect variability in transitions.
@@
Interestingly, while the number of cases in Berlin has little influence on the infection wave in Brandenburg, there is a strong influence in the other direction. This asymmetry is especially visible in our plots when varying the population size, for example, in Figs. 10(a) and 13(a), but especially through the extreme experiments within the same population size. It is important to keep in mind that the population size in Berlin is significantly larger than in Brandenburg and also larger than the total number of agents in the ABM contribution of the hybrid model. On explanation lies in the asymmetry of commuting: according to data from 2021, around 270,000 inhabitants of Brandenburg had their work place in Berlin, while only around 108,000 inhabitants of Berlin had their work place in Brandenburg [12]. These numbers likely underestimate the actual pre-pandemic commuting flows due to increased remote work in 2021. Still, they highlight the significant directional influence of Brandenburg on Berlin.
@@ 7 Conclusion
By applying this hybrid ABM-PDE model to the Berlin-Brandenburg region using real-world mobility and infection data, we demonstrated its ability to capture the distinct epidemiological dynamics of urban and rural areas. The model successfully reduced computational cost significantly compared to a full-ABM simulation. This efficiency gain becomes evident when comparing the simulation times and error margins of the hybrid model and the full-ABM. Specifically, the hybrid model achieved shorter simulation times and lower errors than the ABM, regardless of whether a partial (25%) or full (100%) population sample was used. Thus, our hybrid approach is a practical alternative for real-time epidemiological assessments. A key advantage over full-ABM models such as EpiSim is that our approach allows for efficient simulations of 100% of the population on a standard laptop, making large-scale epidemiological modeling more accessible and computationally feasible.

### v2

@@ 6 Discussion
The deterministic PDE dynamics and the coupling process both reduce variability in the 25% and the 100% scenarios. In fact, the hybrid model introduces a different type of noise due to the abstraction and simplification of agent-level dynamics. Agent identifiers are lost in the PDE region, leading to mismatches when assigning states upon re-entry to the ABM. For instance, an agent may reappear earlier or later than expected, or with an incorrect health status. To mitigate this, we propose future improvements such as storing metadata when an agent transitions into the PDE domain – including expected return time and current health status. Upon re-entry, this information could be used to assign a health state based on time spent in the PDE region and the transition dynamics from the compartment model (see Fig. 1). This strategy would help reduce artificial variance introduced by random reassignment. Additionally, stochasticity could be reintroduced into the PDE model itself – for example, through noise terms in the compartment equations – to better reflect variability in transitions.
@@
Interestingly, while the number of cases in Berlin has little influence on the infection wave in Brandenburg, there is a strong influence in the other direction. This asymmetry is especially visible in our plots when varying the population size, for example, in Figs. 10(a) and 13(a), but especially through the extreme experiments within the same population size. It is important to keep in mind that the population size in Berlin is significantly larger than in Brandenburg and also larger than the total number of agents in the ABM contribution of the hybrid model. One explanation lies in the asymmetry of commuting: according to data from 2021, around 270,000 inhabitants of Brandenburg had their work place in Berlin, while only around 108,000 inhabitants of Berlin had their work place in Brandenburg [11]. These numbers likely underestimate the actual pre-pandemic commuting flows due to increased remote work in 2021. Still, they highlight the significant directional influence of Brandenburg on Berlin.
@@ 7 Conclusion
By applying this hybrid ABM-PDE model to the Berlin-Brandenburg region using real-world mobility and infection data, we demonstrated its ability to capture the distinct epidemiological dynamics of urban and rural areas. The model successfully reduced computational cost significantly compared to a full-ABM simulation. This efficiency gain becomes evident when comparing the simulation times and error margins of the hybrid model and the full-ABM. Specifically, the hybrid model achieved shorter simulation times and lower errors than the ABM, regardless of whether a partial (25%) or full (100%) population sample was used. Thus, our hybrid approach is a practical alternative for real-time epidemiological assessments. A key advantage over full-ABM models such as EpiSim is that our approach allows for efficient simulations of 100% of the population on a standard computer, making large-scale epidemiological modeling more accessible and computationally feasible.